OpenStudy (anonymous):

PLEASE HELP ME!! Solve the following system of equations. 9x + 4y = -45 3x + y = -6

OpenStudy (texaschic101):

do you prefer substitution or elimination ?

OpenStudy (anonymous):

@texaschic101 elimination

OpenStudy (texaschic101):

9x + 4y = -45 3x + y = -6 --->(-4) ------------------ 9x + 4y = -45 -12x - 4y = 24 (result of multiplying by -4) -----------------add -3x = -21 -- divide by 3 x = 7 now sub 7 in for x in either equation to find y 3x + y = -6 3(7) + y = -6 21 + y = -6 y = -6 + - 21 y = -27 check.. 9x + 4y = -45 9(7) + 4(-27) = -45 63 - 108 = -45 -45 = -45 (correct) x = 7 and y = -27

OpenStudy (anonymous):

@texaschic101 thank you so much, and thanks for walking through the problem, i understand it now! thank u

OpenStudy (texaschic101):

your welcome :)